OpenStudy (sriram):

if the locus of point of intersection of tangents to the parabola y^2=4ax which intecept a fixed length L on the directrix is (y^2-Kax) (x+a)^2=(Lx)^2 Then find the value of K.

OpenStudy (sriram):

K is an integer from 1 to 9

OpenStudy (anonymous):

IITJEE integer type?

OpenStudy (sriram):

nope jst asked in my paper

OpenStudy (anonymous):

the tangent is y=mx+a/m the directrix is x=-a let ur desired point be h,k put in h,k into the tangent line k = mh + a/m m^2h -mk +a = 0 find the values of m from this quadratic now solve both the lines with x+a=0 set the difference bgtw the y-coordinates as L